OpenStudy (anonymous):

The area of a triangle is 338 square yards. The height of the triangle is four times its base. Find the height and base of the triangle.

OpenStudy (theopenstudyowl):

This is like an SAT question

OpenStudy (theopenstudyowl):

x=13

OpenStudy (jmark):

Base=x Height=4x 338=(1/2)(x)(4x) 338=4x^2/2 338=2x^2 169=x^2 13=x Base is 13 yards Height is 4*13= 52 yards Follow us at : https://twitter.com/score_more